Wilson bought some flour. He gave 20% of the flour to mike and 25% to Mary. He gave them 72kg of flour in total. He then packed the rest of the flour in to 2 kg and 3 kg bags. There were twice as many 3 kg bags as 2 kg. How many 3 kg bags were there?.
Answer by Theo(13342)

20% to mike and 25% to mary means a total of 45%.
the amount given to mike and mary total 78 kilograms.
if you let x equal the total amount of flour, and the amount given to mike and mary is 45%, then you get .45 * x = 78
solve for x to get x = 78/.45 = 160
160 minus 78 = 88 kilograms that are left after mike and mary got their share.
we can reuse the variable x because it is not needed any more.
if you let x equal the number of 2 kilogram bags and you let y equal the number of 3 kilogram bags, then you get 2x + 3y = 88
you are given that y = 2x, therefore you can replace y in that equation with 2x to get 2x + 2 * 3x = 88.
simplify and combine like terms to get 8x = 88
solve for x to get x = 11
since y = 2x, then you get y = 22
the equation of 2x + 3y = 88 becomes 2 * 11 + 3 * 22 = 88 which becomes 22 + 66 = 88 which becomes 88 = 88, confirming the values for x and y are good.
your question was how many 3 kilogram bags were there.
the answer is 22.